OpenStudy (anonymous):

The function f=9/5c+32 gives the temperature in degrees Fahrenheit, F, based on the temperature in degrees Celsius, C. Use the function to fill the table with the number of degrees Fahrenheit for different numbers of degrees Celsius. Degrees Celsius,c -15 -10 -5 0 5 10 15

OpenStudy (anonymous):

Hi Welcome to Openstudy! You are given the function F=9/5C+32 Only thing you have to do is plug the given values of C into the equation and solve for F. For example, the first one, you are given -15°C. PLug this in F=9/5C+32 it will be like this: F=9/5(-15)+32 get your calculator and solve for F ^_^ Do the same thing with the rest until you complete the table.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome ^_^